OpenStudy (anonymous):

Natasha rides her bike (at a constant speed) for 3 hours, helped by a wind of 3 miles per hour. Pedaling at the same rate, the trip back against the wind takes 9 hours. Find find the total round trip distance she traveled. Natasha rides her bike (at a constant speed) for 3 hours, helped by a wind of 3 miles per hour. Pedaling at the same rate, the trip back against the wind takes 9 hours. Find find the total round trip distance she traveled. @Mathematics

OpenStudy (tastytune):

distance = 27 miles

OpenStudy (anonymous):

r t = d Solve the following for r:\[(r+3)*3=(r-3)*9 \]r = 6 Using the equation's LHS, the round trip distance is\[2*(6+3)*3= 54 \text{ miles} \]

OpenStudy (tastytune):

Yes 27 miles each, therefore 54 miles the roundtrip..good explanation robtobey